Question 2
a li 0. STUDENTS PHY 114: Section. MEDGAR EVERS COLLEGE CITY UNIVERSITY OF NEW YORK DEPARTMENT OF PHYSICS & COMPUTER SCIENCES SPRING 2018: BASIC PHYSICS EXAM#2 03-28-18 Instructor: Dr. Ifeanyi E. Ekejiuba The following formulas may be useful. l. Equations of linear motion: ( 1 ) v Vo + at ; (2) d vst+%ae, (3)v,v/ + 2ad ii. Newtons 2nd law of motion: F ma ili Linear momentum, p- mv AE . Powr, Por P vi. Kinetic Energy, K.E.mv vi. Potential Energy, P.E.-mgh vill. Take: 10 m/s Show your work in the spaces provided( 50 % 1. Then, circle the Letter that corresponds to the correct answer[ 50 % ] INSTRUCTIONS: 1. A 1-kilogram block is pushed across a horizontal surface with a horizontal force of 20 N against a friction force of 1 N. The acceleration of the block is A) 1m/s? B) 2 m/s2 C) 5 mis.D) 10 m/s?. (Eyhone of the above Fa ma An astronaut on another planet drops a 1-kg rock from rest and finds that it falls a vertical distance of 2 meters in one second. On this planet, the rock has a weight of 2. A) 1N B)) 2 N C) 3 N D) 4 N E) 5 N
Answer #1

2) Let the acceleration due to gravity on the planet be g

Then the distance through which the rock falls in time t is

\small d=\frac{gt^2}{2}

Therefore the acceleration due to gravity is

\small g=\frac{2d}{t^2}=\frac{2\times 2}{1^2}=4\,\,m/s^2

The weight of the rock is

\small F=mg=1\times 4 = 4\,N

Anser is option D) 4 N